If 8 is subtracted from a number and the difference is doubled, the results is 5 more than the number. Find the number.
Answer by ana@10(4)

Consider the number to be x. Then by given information by subtracting 8 from x we get certain number..let that number be y. So
x-8=y by rearranging x-y=8
Then by second information double tge difference is 5more than original number..
So 2y=x+5
So by solving both equations we get y=13 amd x=21..
So the number is 21.